OpenStudy (meilendurcer):

The equation d = 16t2 can be used to find the distance, in feet, that objects fall in t seconds. Graph the equation to find how far an object will fall in 2 seconds.

OpenStudy (raden):

just put t=2, so d(2)=16(2)^2 = 16*4 = 64 ft